OpenStudy (anonymous):

what is the solution to the system of equations? x+y=2 x-y=4

OpenStudy (alexandervonhumboldt2):

x=2-y x=4+y 4+y=2-y

OpenStudy (alexandervonhumboldt2):

2y=2 y=1 x=1

OpenStudy (alexandervonhumboldt2):

(1, 1)

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

seems unlikely that \(1-1=4\) right?

OpenStudy (anonymous):

add and get \[2x=6\] then go from there

OpenStudy (anonymous):

(2,6)? @satellite73

OpenStudy (anonymous):

@satellite73